Write the equation of motion and Laplace transform for the following system knowing that Assume zero initial conditions.
Answer #1

given RLC (series RLC) circuit is + rele W 1 t tut Apply KVL :- & - IR-L da - I lidt = 0 IR + L 1 2 3 + L t = Now apply caplasc Iis) E(S) Sict SRC + 1 sc . - LCS + SE + tc ) SIL - S2 + S + TC S112 = - & I J (&) 2 - 1 4 - e 2 51,2 = - = √ (2) ² - to n